When scaling Mt. Everest, climbers usually require supplemental oxygen, because the partial pressure of this gas is too low to meet their O2 demand. What is the partial pressure of O2 in torr at the top of Mt. Everest where the atmospheric pressure is 200 torr, the partial pressure of N2 is 156 torr, and the partial pressure of the other gases (excluding O2) is 2 torr?
1
Expert's answer
2016-07-27T08:16:02-0400
One millimeter of mercury is approximately 1 torr, or 1/760 of standard atmospheric pressure. The two units are not exactly equal; however, the difference (less than 0.000017% to be exact) is negligible for most practical uses.
The partial pressure of the gas mixture is the sum of all gas pressure,
Here of P=P(N2)+p(O2)+P(other)=200(torr)
P(O2)= 200-156-2=42(torr)

Answer: P(O2) is 42torr